OpenStudy (anonymous):

find the slope intercept form for the line satisfying the following conditions x intercept 2 y intercept 3/5

OpenStudy (ivancsc1996):

Slope is\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }=\frac{ 3/5-0 }{ 0-2 }=-\frac{ 3 }{ 10 }=-0.3\]